OpenStudy (anonymous):

Can someone help my factor: 5x^2-21=-4 (I can't seem to get the right numbers..)

hartnn (hartnn):

5x^2-21+4=0 so tell me 2 numbers whose product is 20 and sum is -21

OpenStudy (anonymous):

I want to say 7 x -3 ? but that doesn't make sense...or at least when I tried doing it, it did not fit in correctly...

hartnn (hartnn):

yes, 7,-3 are not correct, the two numbers are simple to find....try again

OpenStudy (anonymous):

I don't know ..

hartnn (hartnn):

try -20 and -1

OpenStudy (anonymous):

WOW....I feel dumb. Thanks for helping me out.

hartnn (hartnn):

welcome....

hartnn (hartnn):

so what factors u got ?

OpenStudy (anonymous):

I plugged in the -21 and -1 : (5x-21)(x-1) and I got 5x^2-5x-21x+21 ..that's wrong isn't it..

hartnn (hartnn):

5x^2-21x+4=0 yeah, coz u need to split the middle term 5x^2-20x-x+4=0

OpenStudy (anonymous):

What? I didn't learn how to split terms...I still don't get how you got that... and how would you factor it?

OpenStudy (anonymous):

wait would the equation become 5x(x-4)(x+4) since you factor out the GCF?

hartnn (hartnn):

5x^2-20x-x+4=0 5x(x-4)-1(x-4)=0 now factor out x-4

OpenStudy (anonymous):

2 and -2 ?

hartnn (hartnn):

5x(x-4)-1(x-4)=0 (x-4)(5x-1)=0 and u are done factoring.....